oh he minimure sal vakue -2n and the maximum unual vake 2 n220.p-04 De not round o-L」(Ibound 50 one decinal pince 8s needed ) μ-2 -□(Round to onedecintel place s needed) +20(Round to one decimal place as needed)
Answer #1

Given n=220 and p=0.4.

\\\mu=np=220(0.4)={\color{Red} 88} \\\\\sigma=\sqrt{np(1-p)}=\sqrt{220(0.4)(1-0.4)}=\sqrt{52.8}={\color{Red} 7.3} \\\\\mu-2\sigma=88-2(7.3)=88-14.6={\color{Red} 73.4} \\\\\mu+2\sigma=88+2(7.3)=88+14.6={\color{Red} 102.6}